You and your business partners are considering applying for a franchise. If approved, you expect startup costs to be $400,000 plus $300,000 in real estate costs. Of this $700,000 in upfront costs, you will be allowed to depreciate $420,000 on a five-year MACRS schedule. The remaining $280,000 is not depreciable but should be included in the book value of the fixed assets associated with the franchise when it is sold. Your plan is to start and operate the business for 5 years at which time you expect to sell the business for $1,000,000. You expect to initially have working capital needs of $30,000, but these needs will grow proportionately with sales. You expect sales in the first year to be $200,000 and that sales will grow by 20% in the second year, 15% in the third year, and then 10% in the fourth and fifth years. You project annual fixed operating expenses of $50,000. Your annual variable operating expenses are expected to be 60% of sales in the first year. With improvements in efficiency and experience, you expect variable operating expenses to be 55% of sales in the second year, 50% of sales in the third, fourth and fifth years. You expect to pay taxes of 20%. Assume your required return is 12%.

3. Estimate the opportunity’s NPV. Explain how you arrived at your NPV estimates in the report.
Answer #1

Book value of the assets after 5 years = BV = Non deprecible base of the upfront cost + remaining book value of the deprecitble base.

Under 5 years MACRS deprecition schedule, pending book value after year 5 = 5.76%

Hence, BV = 280,000 + 420,000 x 5.76% =  304,192

Sale value for the buisness = SV = 1,000,000

Gain on sale, G = SV - BV = 1,000,000 -  304,192 =  695,808

Tax on gain = G x tax rate =  695,808 x 20% =  139,162

Hence, post tax sale value = SV - tax on gain = 1,000,000 - 139,162 =  860,838

Please see the table below. Please be guided by the second column titled “Linkage” to understand the mathematics. The last row highlighted in yellow is your answer. Figures in parenthesis, if any, mean negative values. All financials are in $. Adjacent cells in blue contain the formula in excel I have used to get the final output.

R T U V w 1 2 3 P 54 Year, n Linkage 55 Upfront costs 56 Depreciable base 57 5 years MACRS dep scheduled (700,000) (420,000)

The opportunity's NPV = $  25,253.61

How we arrived at the NPV:

  • As a first step, we identify the relevant cash flows from the project
  • We start with initial investment (cash flows at t = 0), then calculate annual cash flows and then finally the terminal year cash flows.
  • The terminal cash flows usually comprise ofpost tax sale value of the business and release of working capital.
  • Once the relevant, post tax cash flows are in place, we calculated the NPV using the NPV function of excel.